OpenStudy (anonymous):

What does giga mean? 10 + 9. 10 - 9. 10^9. 10^-9.

OpenStudy (anonymous):

will reward all helpers

OpenStudy (welshfella):

giga is 1,000,000,000

OpenStudy (welshfella):

convert that to scientific form

OpenStudy (welshfella):

1,000 = 10^3 10,000 = 10^4 etc

OpenStudy (anonymous):

thank you :D

OpenStudy (welshfella):

yw